OpenStudy (anonymous):

Find the derivative of 7x/9

OpenStudy (anonymous):

Y = log(7x-9) 10^y = 7x-9 Now apply natural log both sides ln(10^y) = ln(7x-9) Yln10 = ln(7x-9) Now differentiate dy/dx*ln(10) = 7/(7x-9) dy/dx =7/{ln(10)*(7x-9)

OpenStudy (anonymous):

Thanks bud =0)

OpenStudy (anonymous):

I'm not sure I understand it, but I'll look it over lata